# Automorphisms of free metabelian Lie algebras

### v2

@@ Abstract.
We show that every automorphism of a free metabelian Lie algebra $M_{n}$ of rank $n\geq 4$ over an arbitrary field $K$ is almost tame, that is, it is a product of so-called Chein automorphisms (or one-row transformations). Moreover, we show that the group of all automorphisms $\mathrm{Aut}(M_{n})$ of $M_{n}$ of rank $n\geq 4$ over a field $K$ of characteristic $\neq 3$ is generated by all linear automorphisms, as well as one quadratic and one cubic automorphism. The same result holds for fields of any characteristic if $n\geq 5$.
We also show that all Chein automorphisms of lower degree $\geq 4$ and all exponential automorphisms of lower degree $\geq 5$ are tame, which contradicts the results of [6, 42].
@@
Mathematics Subject Classification (2020): 17B40, 17B01, 17B30.
Key words: Automorphism, Lie algebra, free metabelian Lie algebra.
@@ 1. Introduction
The well known Jung–van der Kulk Theorem [19, 24] states that every automorphism of the polynomial algebra $K[x,y]$ in two variables $x,y$ over an arbitrary field $K$ is tame. An analogue of this result for free associative algebras in two variables was proven by Makar-Limanov and Czerniakiewicz [15, 28]. The same result is true for free Poisson algebras in two variables over a field of characteristic zero [30] (see also [29, 70, 31]). Moreover, the automorphism groups of polynomial algebras, free associative algebras, and free Poisson algebras in two variables are isomorphic. The tameness of automorphisms of free right-symmetric algebras in two variables over any field was proven in [22].
@@
of the polynomial algebra $K[x,y,z]$ over a field $K$ of characteristic 0 is proven to be wild [53, 66]. The well-known Anick automorphism (see [14, p. 398])
@@
of the free associative algebra $K\langle x,y,z\rangle$ over a field $K$ of characteristic 0 is also proven to be wild [67, 68]. The Nagata automorphism gives an example of a wild automorphism of free Poisson algebras in three variables. In 2006 I.P. Shestakov (unpublished) constructed an analogue of the Anick automorphism
$(x+\{xz-\{y,z\},z\},y+z(xz-\{y,z\}),z)$
for free Poisson algebras in three variables.
@@
The variety of metabelian Lie algebras is also a classical and well-studied variety of Lie algebras. In this paper we denote by $M_{n}$ a free metabelian Lie algebra of rank $n$ over a field $K$. It is well known that any nontrivial exponential automorphism of $M_{2}$ is wild (see, for example [56, Proposition 4],[2, Theorem 3], [44, Theorem 3]). In 1992, Drensky [17] proved that the exponential automorphism $\mathrm{exp}(\mathrm{ad}[x_{1},x_{2}])$ of $M_{3}$ is wild. In 2008, Roman’kov [49] proved that the group of automorphisms of $M_{3}$ cannot be finitely generated modulo the subgroup generated by all exponential and tame automorphisms. Kabanov and Roman’kov [20] constructed strictly nontame primitive elements of $M_{3}$.
In 2015, Nauryzbaev [39] proved that the group of tame automorphisms of $M_{n}$ is generated by all linear automorphisms and the quadratic automorphism
@@
In 1992 V. Drensky [17], in 1993 Papistas [43], and in 1995 Bahturin and Shpilrain [7] proved that free algebras of rank $n\geq 2$ of any polynilpotent variety of Lie algebras that is not abelian or metabelian have wild automorphisms. This result also makes the case of free metabelian Lie algebras more important.
In 1992 Yu. Bahturin and S. Nabiyev [6] claimed that every nontrivial exponential automorphism of the free metabelian Lie algebra $M_{n}$ of any rank $n\geq 2$ over a field of characteristic zero is wild. More examples of wild automorphisms of $M_{n}$ of rank $n\geq 4$ were given in 2008 by Z. Özcurt and N. Ekici [42]. For a long time, the area specialists believed these results were true. Unfortunately, both articles have recently been shown to contain fatal errors [71].
Inner automorphisms of groups play a crucial role in the proofs of Bachmuth-Mochizuki-Roman’kov Theorem [5, 46]. The group automorphism
$(x_{2}^{-1}x_{1}x_{2},x_{2},\ldots,x_{n})$
is extreemly important in their proofs and there is no analogue of this automorphism in the case of free metabelian Lie algebras. This makes the case of algebras more complicated.
One property of the free metabelian Lie algebras is that they admit non-elementary automorphisms that change only one variable and fix the rest. All automorphisms of $M_{n}$ of this type (including elementary ones) are called Chein automorphisms or one-row transformations [18, 48](see definitions in Section 2). The first example of the Chein automorphism is
(2) $(x_{1}+[[x_{2},x_{3}],x_{1}],x_{2},\ldots,x_{n})$
and no one could answer whether it is tame or not. For this reason, the specialists in the area easily accepted the wrong results [6, 42]. This automorphism is wild if $n=3$ [36], and is the first candidate for wildness for all $n\geq 4$.
Chein automorphisms were accepted as almost elementary automorphisms. An automorphism of $M_{n}$ is called almost tame if it is a product of Chein automorphisms [36, 38].
In this paper we show that every automorphism of $M_{n}$ of rank $n\geq 4$ is almost tame. Moreover, $K$ is a field of characteristic $\neq 3$ then the group of all automorphisms $\mathrm{Aut}(M_{n})$ of $M_{n}$ of rank $n\geq 4$ is generated by all linear automorphisms, as well as a quadratic automorphism (1) and a cubic automorphism (2). If $n\geq 5$ then the same result holds for fields of any characteristic.
One of the most important points of the proof is that we notice that all Chein automorphisms of lower degree $\geq 4$ and all exponential automorphisms of lower degree $\geq 5$ are tame. These results contradict the results of [6, 42].
We follow the proof scheme given in [5], and in several places, we repeat the proofs for completeness. The real difficulties arise in those places where inner automorphisms were used in [5].
Thus the question about the existence of wild automorphisms for a free metabelian Lie algebra $M_{n}$ of rank $n\geq 4$ is still open as well as in the case polynomial and free associative algebras. But in this case it is reduced to a concrete automorphism.
Problem 1.
Is the Chein automorphism (2) of the free metabelian algebra Lie algebra $M_{n}$ of rank $n\geq 4$ tame?
The paper is organized as follows. In Section 2 we define some important automorphisms and some subgroups of the automorphism group $\mathrm{Aut}(M_{n})$. In Section 3, using a representation of $M_{n}$ via Magnus embedding, we give the definition of the Fox derivatives and the Jacobian matrix of an endomorphism, and determine some of their properties. Using some representations of automorphisms from this section, in the following sections we make calculations in the language of the Jacobian matrices instead of automorphisms. For this, in Section 4 we clearly determine some matrix notations. In Section 5 we prove that a huge part of Chein and exponential automorphisms are tame. We begin the study of almost tame automorphisms in Section 6. Section 7 is devoted to the study of some special class of matrices that includes the Jacobian matrices of an important class of automorphisms. In Section 8 we prove a lemma on eliminating a variable in some special cases. Section 9 is devoted to the study of one exceptional automorphism
(3) $(x_{1}+[[x_{1},x_{2}],x_{1}],x_{2}+[[x_{1},x_{2}],x_{2}],x_{3},\ldots,x_{n})$
of $M_{n}$. In Section 10 we show how to eliminate a variable in the general case. In Section 11 the main result and some its corollaries are formulated.
2. Automorphisms of $M_{n}$
Let $L_{n}$ be a free Lie algebra over a field $K$ freely generated by $z_{1},z_{2},\ldots,z_{n}$. Then $Y_{n}=L_{n}/[L_{n},L_{n}]$ is an abelian Lie algebra with a linear basis $y_{1},y_{2},\ldots,y_{n}$ and $M_{n}=L_{n}/[[L_{n},L_{n}],[L_{n},L_{n}]]$ is a free metabelian Lie algebra with free generators $x_{1},\ldots,x_{n}$, where $y_{i}=z_{i}+[L_{n},L_{n}]$ and $x_{i}=z_{i}+[[L_{n},L_{n}],[L_{n},L_{n}]]$ for all $1\leq i\leq n$.
Thus $M_{n}^{\prime}=[M_{n},M_{n}]$ is abelian and the set of all right normed (nonassociative) words of the form
is a linear basis of $M_{n}^{\prime}$ (see, for example [8, p. 139]).
Denote by $\phi=(f_{1},\ldots,f_{n})$ the endomorphism of $M_{n}$ such that $\phi(x_{i})=f_{i}$ for all $i$. In particular, $\mathrm{id}=(x_{1},\ldots,x_{n})$ is the identity automorphism. An automorphism of $M_{n}$ of the form
$(x_{1},\ldots,x_{i-1},\alpha x_{i}+f,x_{i+1},\ldots,x_{n}),$
where $0\neq\alpha\in K$ and $f$ does not depend on $x_{i}$, is called elementary. The subgroup $\mathrm{TAut}(M_{n})$ of $\mathrm{Aut}(M_{n})$ generated by all elementary automorphisms is called the tame automorphism group, and the elements of this subgroup are called tame automorphisms of $M_{n}$. Non-tame automorphisms of $M_{n}$ are called wild.
@@ 2. Automorphisms of $M_{n}$
which include all elementary automorphisms but not necessarily elementary ones. They are called Chein automorphisms or one-row transformations [18]. These automorphisms move only one variable and fix the others. The subgroup $\mathrm{ATAut}(M_{n})$ of $\mathrm{Aut}(M_{n})$ generated by all Chein automorphisms will be called the almost tame automorphism group, and the elements of this subgroup will be called called almost tame automorphisms of $M_{n}$ [36]. Thus
@@
If $n=3$ then every non-elementary Chein automorphism is wild [36, 49] and the automorphism (3) is not almost tame [36]. The case $n=3$ is well studied and defining relations for elementary automorphisms [37] and for Chein automorphisms [38] are given.
As usual, the group of all linear automorphisms of $M_{n}$ will be identified with the group of all invertible matrices $GL_{n}(K)$. Recall that the symmetric group $S_{n}$ is embedded into $GL_{n}(K)$ and any permutation $\pi\in S_{n}$ represents the automorphism $\pi=(x_{\pi(1)},\ldots,x_{\pi(n)})$. For example, the transposition $(23)$ represents the automorphism $(23)=(x_{1},x_{3},x_{2},x_{4},\ldots,x_{n})$.
@@
The group of tame automorphisms $\mathrm{TAut}(M_{n})$ is generated by all linear automorphisms $GL_{n}(K)$ and all automorphisms of the form (5) where $f\in M_{n}^{2}$ does not depend on $x_{1}$. Moreover, if $n\geq 4$ and $K$ is a field of characteristic $\neq 3$, then $\mathrm{TAut}(M_{n})$ is generated [39] by all linear automorphisms $GL_{n}(K)$ and the quadratic elementary automorphism (1). If $n\geq 5$ then the same result holds over fields of any characteristic.
Let $0\neq z\in[M_{n},M_{n}]$. Then $\mathrm{ad}\,z:M_{n}\to M_{n}(m\mapsto[z,m])$ is a locally nilpotent derivation and $(\mathrm{ad}\,z)^{2}=0$. The exponential automorphism $\exp(\mathrm{ad}\,z)$ of $M_{n}$ can be written as
$\exp(\mathrm{ad}\,z)=\mathrm{id}+\mathrm{ad}\,z=(y_{1}+[z,y_{1}],\ldots,y_{n}+[z,y_{n}]).$
It is known that the exponential automorphism $\exp(\mathrm{ad}\,[x_{1},x_{2}])$ is wild if $n=3$ [17, 36, 49].
@@
$M_{n}=T_{1}\oplus T_{2}\oplus\ldots\oplus T_{k}\oplus\ldots$
be the standard degree grading of $M_{n}$, i.e., $T_{k}$ is the linear span of (non-associative) monomials in $x_{1},x_{2},\ldots,x_{n}$ of degree $k$. If
$f=f_{s}+f_{s+1}+\ldots+f_{t}\in M_{n},f_{j}\in T_{j},f_{s}\neq 0,f_{t}\neq 0,$
then $\deg(f)=t$ is the degree and $\mathrm{ldeg}(f)=s$ is the lower degree of $f$ in the variables $x_{1},x_{2},\ldots,x_{n}$. Set also $\deg(0)=$-$\infty$ and $\mathrm{ldeg}(0)=\infty$.
Let
$\phi=(x_{1}+f_{1},x_{2}+f_{2},\ldots,x_{n}+f_{n}),$
where $f_{i}\in M_{n}^{2}$ for all $1\leq i\leq n$. Then we define $\deg(\phi)$ as the maximum of all $\deg(f_{i}),1\leq i\leq n$, and $\mathrm{ldeg}(\phi)$ as the minimum of all $\mathrm{ldeg}(f_{i}),1\leq i\leq n$. If $\deg(\phi)=\mathrm{ldeg}(\phi)=k\geq 1$ then we say that $\phi$ is homogeneous of degree $k$. In particular, if $k=2$ then $\phi$ is called quadratic and if $k=3$ then it is called cubic.
The group of automorphisms $\mathrm{IAut}(M_{n})$ of $M_{n}$ that induces the identity automorphism on $M_{n}/[M_{n},M_{n}]$ plays an important role in our study. Obviously, $\mathrm{IAut}(M_{n})$ is the set of all automorphisms of $M_{n}$ of lower degree $\geq 2$. The elements of $\mathrm{IAut}(M_{n})$ are called $\mathrm{IA}$-automorphisms.
Here we describe a set of generators of the group of tame $\mathrm{IA}$-automorphisms and almost tame $\mathrm{IA}$-automorphisms.
Lemma 1.
$(a)$ The group of tame $\mathrm{IA}$-automorphisms $\mathrm{TAut}(M_{n})\cap\mathrm{IAut}(M_{n})$ is generated by all linear conjugates of elementary automorphisms of the form (5) where $f\in[M_{n},M_{n}]$ does not depend on $x_{1}$.
$(b)$ The group almost tame $\mathrm{IA}$-automorphisms $\mathrm{TAut}(M_{n})\cap\mathrm{IAut}(M_{n})$ is generated by all linear conjugates of Chein automorphisms of the form (5).
Proof. Every tame automorphism $\phi$ can be represented as
$\phi=\alpha_{1}\phi_{1}\alpha_{2}\phi_{2}\ldots\phi_{k}\alpha_{k+1},$
where $\phi_{i}$ is an elementary automorphism of the form (5) and $\alpha_{i}$ is linear for all $i$. Since $\phi\in\mathrm{IAut}(M_{n})$ for all $i$ it follows that
$\alpha_{1}\alpha_{2}\ldots a_{k+1}=\mathrm{id}.$
since $\phi_{i}\in\mathrm{IAut}(M_{n})$ for all $i$. Then $\alpha_{k+1}=(\alpha_{1}\ldots\alpha_{k})^{-1}$ and, consequently,
$\phi=(\alpha_{1}\phi_{1}\alpha_{1}^{-1})(\alpha_{1}\alpha_{2}\phi_{2}(\alpha_{1}\alpha_{2})^{-1})\ldots(\alpha_{1}\ldots\alpha_{k}\phi_{k}(\alpha_{1}\ldots\alpha_{k})^{-1})$
The statement $(b)$ of the lemma can be proved similarly. $\Box$
3. Jacobian matrices
In this section, we define and state some necessary information about Fox derivatives and Jacobian matrices from [45, 58, 62, 64]. In fact, we define Fox derivatives more directly for free metabelian Lie algebras.
First we give another representation of the free metabelian algebra $M_{n}$ over $K$ with free generators $x_{1},\ldots,x_{n}$. Let $Y_{n}$ be the abelian Lie algebra with a linear basis $y_{1},y_{2},\ldots,y_{n}$. The universal enveloping algebra $U(Y_{n})$ of $Y_{n}$ is the polynomial algebra $U=U(Y_{n})=K[y_{1},\ldots,y_{n}]$ in the variables $y_{1},y_{2},\ldots,y_{n}$. Let

### v3

@@ Abstract.
We show that all Chein automorphisms (or one-row transformations) of lower degree $\geq 4$ of a free metabelian Lie algebra $M_{n}$ of rank $n\geq 4$ over an arbitrary field $K$ of characteristic $\neq 3$ are tame. We then show that all exponential automorphisms of $M_{n}$ of lower degree $\geq 5$ are also tame under the same conditions. The same results hold for fields of any characteristic when $n\geq 5$. These results contradict some long-standing results in the area.
We also prove that a large class of automorphisms of $M_{n}$ of rank $n\geq 4$ that move only two variables are almost tame, that is, they can be expressed as a product of Chein automorphisms.
@@
Mathematics Subject Classification (2020): 17B40, 17B01, 17B30, 16W20.
Key words: Automorphism, derivation, free metabelian Lie algebra.
@@ 1. Introduction
The well known Jung–van der Kulk Theorem [16, 20] states that every automorphism of the polynomial algebra $K[x,y]$ in two variables $x,y$ over an arbitrary field $K$ is tame. An analogue of this result for free associative algebras in two variables was proven by Czerniakiewicz and Makar-Limanov [12, 24]. The same result is true for free Poisson algebras in two variables over a field of characteristic zero [26] (see also [25, 27, 65]). Moreover, the automorphism groups of polynomial algebras, free associative algebras, and free Poisson algebras in two variables are isomorphic. The tameness of automorphisms of free right-symmetric algebras in two variables was proven in [18].
@@
of the polynomial algebra $K[x,y,z]$ over a field $K$ of characteristic 0 is proven to be wild [46, 60]. The well-known Anick automorphism (see [11, p. 398])
@@
of the free associative algebra $K\langle x,y,z\rangle$ over a field $K$ of characteristic 0 is also proven to be wild [61, 63]. The Nagata automorphism gives an example of a wild automorphism of free Poisson algebras in three variables. Recently Shestakov and Zhang [48] constructed an analogue of the Anick automorphism for free Poisson algebras in three variables.
@@
The variety of metabelian Lie algebras is also a classical and well-studied variety of Lie algebras. Throughout this paper, $M_{n}$ denotes the free metabelian Lie algebra of rank $n$ over a field $K$ generated by $x_{1},\ldots,x_{n}$. It is well known that any nontrivial exponential automorphism of $M_{2}$ is wild (see, for example [1, 51]). In 1992, Drensky [14] proved that the exponential automorphism $\mathrm{exp}(\mathrm{ad}[x_{1},x_{2}])$ of $M_{3}$ is wild. More examples of wild automorphisms of $M_{3}$ are given in [17, 32, 43].
In 1992 Bahturin and Nabiyev [5] claimed that every nontrivial exponential automorphism of the free metabelian Lie algebra $M_{n}$ of any rank $n\geq 2$ over a field of characteristic zero is wild. More examples of wild automorphisms of $M_{n}$ of rank $n\geq 4$ were given in 2008 by Özcurt and Ekici [39]. For a long time, the area specialists believed these results were true. Unfortunately, both articles have fatal errors [66].
We show that all Chein automorphisms (or one-row transformations) of lower degree $\geq 4$ of a free metabelian Lie algebra $M_{n}$ of rank $n\geq 4$ over an arbitrary field $K$ of characteristic $\neq 3$ are tame. We then show that all exponential automorphisms of $M_{n}$ of lower degree $\geq 5$ are also tame under the same conditions. The same results hold for fields of any characteristic when $n\geq 5$. These results contradict those of [5, 39].
We also prove that a large important class of automorphisms of $M_{n}$ of rank $n\geq 4$ that move only two variables are almost tame, that is, they can be expressed as a product of Chein automorphisms.
More precise formulations of the results are given in the following sections.
The paper is organized as follows. In Section 2 we define some important automorphisms of $M_{n}$ and give precise formulations some of our results. In Section 3, using a representation of $M_{n}$ via Magnus embedding and Fox derivatives, we describe Chein autoimorphisms of $M_{n}$. In Section 4 we prove that a huge part of Chein and exponential automorphisms are tame. In Section 5 we prove that a huge class of automorphisms that move only two variables are almost tame.
2. Automorphisms of $M_{n}$
Let $M_{n}$ be the free metabelian Lie algebra over a field $K$ with free generators $x_{1},\ldots,x_{n}$. The set of all right normed Lie monomials of the form
is a linear basis of $M_{n}^{\prime}=[M_{n},M_{n}]$ (see, for example [6, p. 139]).
Let $\mathrm{Aut}(M_{n})$ be the group of all automorphisms of $M_{n}$. An automorphism of $M_{n}$ of the form
$(x_{1},x_{2},\ldots,x_{i-1},\alpha x_{i}+f,x_{i+1},\ldots x_{n}),$
where $0\neq\alpha\in K$ and $f$ does not contain $x_{i}$, is called elementary. The subgroup $\mathrm{TAut}(M_{n})$ of $\mathrm{Aut}(M_{n})$ generated by all elementary automorphisms is called the tame automorphism group, and the elements of this subgroup are called tame automorphisms of $M_{n}$. Nontame automorphisms of $M_{n}$ are called wild.
Defining relations among elementary automorphisms have been described for polynomial algebras in three variables [61, 62], for finitely generated free Lie algebras [64], and for free metabelian Lie algebras in three variables [33].
As usual, the group of all linear automorphisms of $M_{n}$ will be identified with the group of all invertible matrices $GL_{n}(K)$. For every pair of integers $s,t$ with $1\leq s\neq t\leq n$, let $(st)$ denote the automorphism that simply permutes the variables $x_{s}$ and $x_{t}$ and leaves all other variables unchanged. The group generated by all $(st)$ is isomorphic to the symmetric group $S_{n}$.
In 2015, Nauryzbaev [35] proved that the group of tame automorphisms of $M_{n}$ is generated by all linear automorphisms $GL_{n}(K)$ and the quadratic automorphism
@@ 2. Automorphisms of $M_{n}$
If $\phi=(f_{1},\ldots,f_{n}),\psi=(g_{1},\ldots,g_{n})\in\mathrm{Aut}(M_{n})$, then the composition $\phi\psi$ is defined as
$\phi\psi=(g_{1}(f_{1},\ldots,f_{n}),\ldots,g_{n}(f_{1},\ldots,f_{n})).$
We also adopt the standard notation $\phi,\psi]=\phi\psi\phi^{-1}\psi^{-1}$ for the commutator, and $\phi^{\psi}=\psi\phi\psi^{-1}$ for conjugation.
Let $0\neq z\in[M_{n},M_{n}]$. Then $\mathrm{ad}\,z:M_{n}\to M_{n}(m\mapsto[z,m])$ is a nilpotent derivation and $(\mathrm{ad}\,z)^{2}=0$. The exponential automorphism $\exp(\mathrm{ad}\,z)$ of $M_{n}$ can be written as
$\exp(\mathrm{ad}\,z)=\mathrm{id}+\mathrm{ad}\,z=(x_{1}+[z,x_{1}],\ldots,x_{n}+[z,x_{n}]).$
It is known that the exponential automorphism $\exp(\mathrm{ad}\,[x_{1},x_{2}])$ of $M_{3}$ is wild [14].
Recall that in 1993 Bryant and Drensky proved [7] that every automorphism of $M_{n}$ of rank $n\geq 4$ over a field $K$ of characteristic zero is approximately tame ( see the definitions in [47]); that is, it can be approximated by tame automorphisms with respect to power series topology [7, 47]. In contrast, all wild automorphisms of $M_{3}$ are absolutely wild [47], meaning that they cannot be approximated by tame automorphisms.
@@
which include all elementary automorphisms but not necessarily elementary ones. They are called Chein automorphisms or one-row transformations [15, 42]. These automorphisms move only one variable and fix the others. The subgroup $\mathrm{ATAut}(M_{n})$ of $\mathrm{Aut}(M_{n})$ generated by all Chein automorphisms is called the almost tame automorphism group, and the elements of this subgroup is called almost tame automorphisms of $M_{n}$ [32, 34]. Thus
@@
The lowest-degree example of a non-elementary Chein automorphism is
(4) $(x_{1}+[[x_{2},x_{3}],x_{1}],x_{2},\ldots,x_{n}).$
This automorphism is wild if $n=3$ [32, 43], and is the first candidate for wildness for all $n\geq 4$, but it is still unknown whether it is tame or not. It was recently shown [36] that an analogue of this automorphism for free metabelian anticommutative algebras is absolutely wild for all $n\geq 3$.
In 2008, Roman’kov [43] proved that the group of automorphisms of $M_{3}$ cannot be finitely generated modulo the subgroup generated by all exponential and Chein automorphisms. Kabanov and Roman’kov [17] constructed strictly nontame primitive elements of $M_{3}$. Defining relations for Chein automorphisms of $M_{3}$ are given in [34].
We introduce the notion of the lower degree of an automorphism. Consider the standard degree grading of
$M_{n}=L_{1}\oplus L_{2}\oplus\ldots\oplus L_{k}\oplus\ldots,$
where $L_{k}$ is the linear span of all Lie monomials of degree $k$ in $x_{1},\ldots,x_{n}$. Every nonzero $m\in M_{n}$ can be uniquely written in the form
$m=l_{s}+l_{s+1}+\ldots+l_{t},$
where $l_{j}\in L_{j}$ for all $s\leq j\leq t$, $l_{s}\neq 0$, and $l_{t}\neq 0$. We call $\mathrm{ldeg}(m)=s$ the lower degree of $m$, while $\deg(m)=t$ denotes the degree of $m$. If $\mathrm{ldeg}(m)=\deg(m)=t$, then $m$ is homogeneous of degree $t$. Set also $\mathrm{ldeg}(0)=\infty$ and $\deg(0)=$-$\infty$.
@@
$\phi=(l_{1}+m_{1},\ldots,l_{n}+m_{n})$
be an automorphism of $M_{n}$ such that $l_{i}\in L_{1}$ and $m_{i}\in[M_{n},M_{n}]$ for all $i$. Set
$\mathrm{ldeg}(\phi)=\mathrm{min}\{\mathrm{ldeg}(m_{i})|1\leq i\leq n\}$
and
$\mathrm{deg}(\phi)=\mathrm{max}\{\mathrm{deg}(m_{i})|1\leq i\leq n\}.$
If $\mathrm{ldeg}(\phi)=\mathrm{deg}(\phi)=t$, then we say that $\phi$ is homogeneous of degree $t$. In particular, (2) is quadratic and (4) is cubic. Clearly, every nonlinear automorphism has a finite lower degree and degree.
We show that all Chein automorphisms of lower degree $\geq 4$ are tame. This means that all Chein automorphisms ”except” (4) are tame. We also show that all exponential automorphisms of lower degree $\geq 5$ are tame. These results contradict the results of [5, 39].
3. Chein automorphisms
In this section, using an analogue of the well known Magnus embedding [19, 23], we give another representation of the free metabelian algebra $M_{n}$, and give direct definition of the Fox derivatives for $M_{n}$. This construction for Lie algebras is a particular case of Shmel’kin’s wreath products [51].
Let $Y_{n}=M_{n}/[M_{n},M_{n}]$ be the abelian Lie algebra with a linear basis $y_{1},y_{2},\ldots,y_{n}$, where $y_{i}=x_{i}+[M_{n},M_{n}]$. The universal enveloping algebra $U(Y_{n})$ of $Y_{n}$ is the polynomial algebra $U=U(Y_{n})=K[y_{1},\ldots,y_{n}]$ in the variables $y_{1},y_{2},\ldots,y_{n}$. Let